OpenStudy (anonymous):

Solve 4x - 2x - 2 = -42.

OpenStudy (anonymous):

4x - 2x -2 = -42 4x-2x=2x 2x - 2 = -42 add 2 to both sides 2x = -40 divide by 2 x = -20